OpenStudy (josee):

Find an equation for the tangent line at the point P (−2, f(−2)) on the graph of f(x) = 3x^2 + 2x + 4

OpenStudy (anonymous):

f(-2)=3(-2)^2+2(-2)+4=12 point P(-2,12)

OpenStudy (josee):

1. y − 10x − 8 = 0 2. y + 10x + 16 = 0 3. y - 10x + 16 = 0 4. y + 10x + 8 = 0 5. y + 10x - 16 = 0 6. y - 10x + 8 = 0 These are the answer choices

OpenStudy (anonymous):

you draw the line of point (-2,12) and then find the slopes and then find the formula